# REvolve: Reward Evolution with Large Language Models for Autonomous Driving

### v1

@@ Appendix D Prompts
1. You may find it helpful to normalize the reward to a fixed reward range like -1,1 or 0,1 by applying transformations like tf.exp() to the overall reward or its components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and not an input variable. Each transformed reward component should have its own temperature variable, and you must carefully set its value based on its effect on the overall reward. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable.
Appendix E Human Feedback Examples
Box 7: REvolve Mutation Prompt Human feedback on the autonomous driving agent’s performance: Positive: collision avoidance, lane keeping, managing consistent speed, intersection and turn handling.
Negative: Smooth Steering.
Based on the human feedback and analysis of the reward components, we can identify that the "smooth driving bonus" needs modification. The feedback indicates issues with smooth steering, which suggests that the existing smoothness component isn’t effectively promoting the desired steering behavior. Given that the human feedback specifically pointed out deficiencies in this area, it suggests the measure (variance of steering actions) might not be adequately sensitive or is scaled improperly.
Intended Mutation: The mutation will focus on revising the "smooth driving bonus" to better capture and reward smooth steering. The current method uses the standard deviation of steering actions to compute a penalty, which may not effectively discriminate between slightly varying and highly erratic steering actions. A more responsive approach might involve squaring the deviation before scaling, which penalizes larger variations more severely and provides a sharper gradient for the learning algorithm to optimize against.
Rationale By squaring the deviation, small steering adjustments are less penalized, allowing for natural minor corrections while driving. Larger, more erratic movements result in significantly higher penalties, pushing the training to favor smoother transitions and more stable steering behaviors.
def old_compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: ... max_steering_variance = 0.1 smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor / max_steering_variance)) ... def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: ... # Updated mutation for smooth driving bonus max_steering_variance = 0.1 smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor**2 / max_steering_variance**2)) ...
Appendix F Additional Results
In Box F we showcase the best performing reward (fitness score of $0.84$) In Box F we showcase the best performing reward (fitness score of $0.81$)
@@ Appendix F Additional Results
Box 8: Best reward function from REvolve def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, min_pos: float, action_list: List[float]) -> Tuple[float, Dict[str, float]]: # Parameters to tweak the importance of different reward components collision_penalty = -100 inactivity_penalty = -10 speed_reward_weight = 2.0 position_reward_weight = 1.0 smoothness_reward_weight = 0.5 # Adjusted temperature parameters for score transformation speed_temp = 0.5 # Increased temperature for speed_reward position_temp = 0.1 smoothness_temp = 0.1 reward_components = { "collision_penalty": 0, "inactivity_penalty": 0, "speed_reward": 0, "position_reward": 0, "smoothness_reward": 0 } # Penalize for collision if collision: reward_components["collision_penalty"] = collision_penalty # Penalize for inactivity (speed too close to zero) if speed < 4.5: # Adjusted threshold for inactivity reward_components["inactivity_penalty"] = inactivity_penalty # Reward for maintaining an optimal speed range if 9.0 <= speed <= 10.5: speed_score = 1 - np.abs(speed - 9.75) / 1.75 # Centering and normalizing around the average of optimal speed range else: speed_score = -1 # Use a sigmoid function to smoothly transform the speed score and constrain it reward_components["speed_reward"] = speed_reward_weight * (1 / (1 + np.exp(-speed_score / speed_temp))) # Reward for being close to the center of the road position_score = np.exp(-min_pos / position_temp) reward_components["position_reward"] = position_reward_weight * position_score # Reward for smooth driving (small variations in consecutive steering actions) steering_smoothness = -np.std(action_list) reward_components["smoothness_reward"] = smoothness_reward_weight * np.exp(steering_smoothness / smoothness_temp) # Calculate total reward, giving precedence to penalties total_reward = 0 if reward_components["collision_penalty"] < 0: total_reward = reward_components["collision_penalty"] elif reward_components["inactivity_penalty"] < 0: total_reward = reward_components["inactivity_penalty"] else: total_reward = sum(reward_components.values()) # Ensure the total reward is within a reasonable range total_reward = np.clip(total_reward, -1, 1) return total_reward, reward_components
@@
Box 9: Best reward function from REvolve Auto def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: # Define temperature parameters for reward transformations temp_collision = 5.0 temp_inactivity = 10.0 temp_centering = 2.0 temp_speed = 2.0 temp_smoothness = 1.0 temp_proximity = 2.0 # Initialize the total reward and reward components dictionary reward_components = { ’collision_penalty’: 0, ’inactivity_penalty’: 0, ’lane_centering_bonus’: 0, ’speed_regulation_bonus’: 0, ’smooth_driving_bonus’: 0, ’front_distance_bonus’: 0 } # Penalty for collisions reward_components[’collision_penalty’] = np.exp(-temp_collision) if collision else 0 # Penalty for inactivity inactivity_threshold = 4.5 if speed < inactivity_threshold and not collision: reward_components[’inactivity_penalty’] = np.exp(-temp_inactivity * (inactivity_threshold - speed)) # Lane centering bonus max_distance_from_center = 2.0 # adaptable based on road width reward_components[’lane_centering_bonus’] = np.exp(-temp_centering * min_pos) # Speed regulation bonus ideal_speed = (9.0 + 10.5) / 2 speed_range = 10.5 - 9.0 reward_components[’speed_regulation_bonus’] = np.exp(-temp_speed * (abs(speed - ideal_speed) / speed_range)) # Smooth driving bonus max_steering_variance = 0.1 # configured to the acceptable steering variance for full bonus smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor / max_steering_variance)) # Front distance bonus max_distance_view = 20 reward_components[’front_distance_bonus’] = np.clip(distance / max_distance_view, 0, 1) # Sum the individual rewards for the total reward total_reward = sum(reward_components.values()) # Ensure penalties take precedence if collision or speed < inactivity_threshold: total_reward = -1 # Apply max negative reward for collision or inactivity # Normalize the total reward to lie between -1 and 1 total_reward = np.clip(total_reward, -1, 1) return total_reward, reward_components
Figure 9: Compares REvolve with baselines using manually designed fitness scores (average score across populations). Again REvolve consistently improves over generations, achieving a higher average fitness score $\sigma=0.67$.

### v2

@@ Appendix E Prompts
The code output should be formatted as a Python code string: ```python``` Some helpful tips for writing the reward function code: 1. You may find it helpful to normalize the reward to a fixed reward range like -1,1 or 0,1 by applying transformations like tf.exp() to the overall reward or its components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and not an input variable. Each transformed reward component should have its own temperature variable, and you must carefully set its value based on its effect on the overall reward. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable.
Box 7: System Prompt: Humanoid Locomotion You are a reward engineer trying to write a reward function for humanoid robot to run. Your goal is to write a reward function for the environment that will help the agent learn the task described in text. Task description: The task is to train a humanoid agent to run. The robot becomes unhealthy if its torso’s z-position falls outside the 1.0 to 2.0 range. The episode ends if the robot becomes unhealthy or after 1000 timesteps, with success measured by speed. Agent locomotion should be more human like. Your reward function should use useful variables from the environment as inputs. An example reward function signature can be: ```python def compute_reward(object_pos: type, goal_pos: type) -> Tuple[type, Dict[str, type]]: ... return reward, {} ``` The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component. The code output should be formatted as a Python code string: ”“‘python … “‘”. 1. You may find it helpful to normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward and/or its reward components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable and assume variables that are not available. 5. Your output must always be the def function code with the output being the total reward. 6. Make sure you don’t give contradictory reward components.
Box 8: Abstracted Environment Variables: Humanoid Locomotion (Part 1) class Env(utils.EzPickle): ## Action Space The action space is a Box(-1, 1, (17,), float32)‘. An action represents the torques applied at the hinge joints. ## Observation Space Observations consist of positional values of different body parts of the Humanoid, followed by the velocities of those individual parts (their derivatives) with all the positions ordered before all the velocities. By default, observations do not include the x- and y-coordinates of the torso. These may be included by passing exclude_current_positions_from_observation=False during construction. In that case, the observation space will be a Box(-Inf, Inf, (378,), float64)‘ where the first two observations represent the x- and y-coordinates of the torso. Regardless of whether exclude_current_positions_from_observation was set to true or false, the x- and y-coordinates will be returned in info with keys "x_position" and "y_position"‘, respectively. However, by default, the observation is a Box(-Inf, Inf, (376,), float64)‘. The elements correspond to the following: | Num | Observation | Min | Max | Name (in corresponding XML file) | Joint | Unit | ---------- | ---- | --- | -------------------------------- | ----- | -------------------------- | | 0 | z-coordinate of the torso (centre) | -Inf | Inf | root | free | position (m) | | 1 | x-orientation of the torso (centre) | -Inf | Inf | root | free | angle (rad) | | 2 | y-orientation of the torso (centre) | -Inf | Inf | root | free | angle (rad) | | 3 | z-orientation of the torso (centre) | -Inf | Inf | root | free | angle (rad) | ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... | 43 | coordinate-2 (multi-axis) of the angular velocity of the angle between torso and left arm (in left_upper_arm) | -Inf | Inf | left_shoulder2 | hinge | anglular velocity (rad/s) | | 44 | angular velocity of the angle between left upper arm and left_lower_arm | -Inf | Inf | left_elbow | hinge | anglular velocity (rad/s) | | excluded | x-coordinate of the torso (centre) | -Inf | Inf | root | free | position (m) | | excluded | y-coordinate of the torso (centre) | -Inf | Inf | root | free | position (m) | Additionally, after all the positional and velocity based values in the table, the observation contains (in order): - *cinert:* Mass and inertia of a single rigid body relative to the center of mass (this is an intermediate result of transition). It has shape 14*10 (*nbody * 10*) and hence adds to another 140 elements in the state space. - *cvel:* Center of mass based velocity. It has shape 14 * 6 (*nbody * 6*) and hence adds another 84 elements in the state space - *qfrc_actuator:* Constraint force generated as the actuator force. This has shape ‘(23,)‘ *(nv * 1)* and hence adds another 23 elements to the state space. - *cfrc_ext:* This is the center of mass based external force on the body. It has shape 14 * 6 (*nbody * 6*) and hence adds to another 84 elements in the state space. where *nbody* stands for the number of bodies in the robot and *nv* stands for the number of degrees of freedom (*= dim(qvel)*)
Abstracted Environment Variables: Humanoid Lomoton (Part 2) def _get_obs(self): device = torch.device("cuda" if torch.cuda.is_available() else "cpu") position = self.data.qpos.flat.copy() velocity = self.data.qvel.flat.copy() com_inertia = self.data.cinert.flat.copy() com_velocity = self.data.cvel.flat.copy() actuator_forces = self.data.qfrc_actuator.flat.copy() external_contact_forces = self.data.cfrc_ext.flat.copy() observation = np.concatenate( (position, velocity, com_inertia, com_velocity, actuator_forces, external_contact_forces,) ) return observation
Box 10: System Prompt: Adroit Hand You are a reward engineer trying to write a reward function for the environment that will help the agent learn the task described in text. Task description: The task involves using an anthropomorphic robotic hand to rotate a door handle and open the door as fast as possible. The robot must manipulate the handle effectively, overcoming the friction and resistance of the door mechanism. The episode ends after 400 steps or if it opened the door successfully. Your reward function should use useful variables from the environment as inputs. An example reward function signature can be: ```python def compute_reward(object_pos: type, goal_pos: type) -> Tuple[type, Dict[str, type]]: ... return reward, {} ``` The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component. The code output should be formatted as a Python code string: ”“‘python … “‘”. 1. You may find it helpful to normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward and/or its reward components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable and assume variables that are not available. 5. Your output must always be the def function code with the output being the total reward. 6. Make sure you don’t give contradictory reward components.
Box 11: Abstracted Environment Variables: Adroit Hand class Env(EzPickle): ## Action Space The action space is a Box(-1.0, 1.0, (28,), float32)‘. The control actions are absolute angular positions of the Adroit hand joints. The input of the control actions is set to a range between -1 and 1 by scaling the real actuator angle ranges in radians. ## Observation Space The observation space is of the type Box(-inf, inf, (97,), float64)‘. It contains information about the angular position of the finger joints, the pose of the palm of the hand, as well as state of the latch and door and the joint velocities and forces. | Num | Observation | Min | Max | Joint Name (in corresponding XML file) | Site Name (in corresponding XML file) | Joint Type| Unit |-----|---------- -|--------|--------|------------| | 0 | Angular position of the vertical arm joint | -Inf | Inf | ARRx | - | hinge | angle (rad) | | ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... ... def _get_obs(self): # qpos for hand # xpos for obj # xpos for target qpos = self.data.qpos.ravel() handle_pos = self.data.site_xpos[self.handle_site_id].ravel() palm_pos = self.data.site_xpos[self.grasp_site_id].ravel() door_pos = np.array([self.data.qpos[self.door_hinge_addrs]]) joint_velocities = self.data.qvel.ravel() joint_forces = self.data.actuator_force.ravel() if door_pos > 1.0: door_open = 1.0 else: door_open = -1.0 latch_pos = qpos[-1] return np.concatenate( [ qpos[1:-2], [latch_pos], door_pos, palm_pos, palm_pos - handle_pos, [door_open] )
Appendix F Human Feedback Examples
Box 12: Mutation with Human Feedback Human feedback on the autonomous driving agent’s performance: Positive: collision avoidance, lane keeping, managing consistent speed, intersection and turn handling. Negative: Smooth Steering. Based on the human feedback and analysis of the reward components, we can identify that the ”smooth driving bonus” needs modification. The feedback indicates issues with smooth steering, which suggests that the existing smoothness component isn’t effectively promoting the desired steering behavior. Given that the human feedback specifically pointed out deficiencies in this area, it suggests the measure (variance of steering actions) might not be adequately sensitive or is scaled improperly. Intended Mutation: The mutation will focus on revising the ”smooth driving bonus” to better capture and reward smooth steering. The current method uses the standard deviation of steering actions to compute a penalty, which may not effectively discriminate between slightly varying and highly erratic steering actions. A more responsive approach might involve squaring the deviation before scaling, which penalizes larger variations more severely and provides a sharper gradient for the learning algorithm to optimize against. Rationale By squaring the deviation, small steering adjustments are less penalized, allowing for natural minor corrections while driving. Larger, more erratic movements result in significantly higher penalties, pushing the training to favor smoother transitions and more stable steering behaviors. def old_compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: ... max_steering_variance = 0.1 smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor / max_steering_variance)) ... def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: ... # Updated mutation for smooth driving bonus max_steering_variance = 0.1 smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor**2 / max_steering_variance**2)) ...
Appendix G Additional Results
G.1 Best performing rewards
G.1.1 Autonomous Driving
In Box G.1.1, we present the best-performing reward from REvolve with a fitness score of $0.84$ for the AirSim environment. Similarly, Box G.1.1 displays the best-performing reward from REvolve Auto with a fitness score of $0.81$ for the same environment.
@@ G.1.1 Autonomous Driving
Box 13: Best reward function from REvolve: Autonomous Driving def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, min_pos: float, action_list: List[float]) -> Tuple[float, Dict[str, float]]: # Parameters to tweak the importance of different reward components collision_penalty = -100 inactivity_penalty = -10 speed_reward_weight = 2.0 position_reward_weight = 1.0 smoothness_reward_weight = 0.5 # Adjusted temperature parameters for score transformation speed_temp = 0.5 # Increased temperature for speed_reward position_temp = 0.1 smoothness_temp = 0.1 reward_components = { "collision_penalty": 0, "inactivity_penalty": 0, "speed_reward": 0, "position_reward": 0, "smoothness_reward": 0 } # Penalize for collision if collision: reward_components["collision_penalty"] = collision_penalty # Penalize for inactivity (speed too close to zero) if speed < 4.5: # Adjusted threshold for inactivity reward_components["inactivity_penalty"] = inactivity_penalty # Reward for maintaining an optimal speed range if 9.0 <= speed <= 10.5: speed_score = 1 - np.abs(speed - 9.75) / 1.75 # Centering and normalizing around the average of optimal speed range else: speed_score = -1 # Use a sigmoid function to smoothly transform the speed score and constrain it reward_components["speed_reward"] = speed_reward_weight * (1 / (1 + np.exp(-speed_score / speed_temp))) # Reward for being close to the center of the road position_score = np.exp(-min_pos / position_temp) reward_components["position_reward"] = position_reward_weight * position_score # Reward for smooth driving (small variations in consecutive steering actions) steering_smoothness = -np.std(action_list) reward_components["smoothness_reward"] = smoothness_reward_weight * np.exp(steering_smoothness / smoothness_temp) # Calculate total reward, giving precedence to penalties total_reward = 0 if reward_components["collision_penalty"] < 0: total_reward = reward_components["collision_penalty"] elif reward_components["inactivity_penalty"] < 0: total_reward = reward_components["inactivity_penalty"] else: total_reward = sum(reward_components.values()) # Ensure the total reward is within a reasonable range total_reward = np.clip(total_reward, -1, 1) return total_reward, reward_components
@@
Box 14: Best reward function from REvolve Auto: Autonomous Driving def compute_reward(curr_x: float, curr_y: float, speed: float, collision: bool, action_list: list, total_step_counter: int, episode_step_counter: int, min_pos: float, distance: float) -> Tuple[float, Dict[str, float]]: # Define temperature parameters for reward transformations temp_collision = 5.0 temp_inactivity = 10.0 temp_centering = 2.0 temp_speed = 2.0 temp_smoothness = 1.0 temp_proximity = 2.0 # Initialize the total reward and reward components dictionary reward_components = { ’collision_penalty’: 0, ’inactivity_penalty’: 0, ’lane_centering_bonus’: 0, ’speed_regulation_bonus’: 0, ’smooth_driving_bonus’: 0, ’front_distance_bonus’: 0 } # Penalty for collisions reward_components[’collision_penalty’] = np.exp(-temp_collision) if collision else 0 # Penalty for inactivity inactivity_threshold = 4.5 if speed < inactivity_threshold and not collision: reward_components[’inactivity_penalty’] = np.exp(-temp_inactivity * (inactivity_threshold - speed)) # Lane centering bonus max_distance_from_center = 2.0 # adaptable based on road width reward_components[’lane_centering_bonus’] = np.exp(-temp_centering * min_pos) # Speed regulation bonus ideal_speed = (9.0 + 10.5) / 2 speed_range = 10.5 - 9.0 reward_components[’speed_regulation_bonus’] = np.exp(-temp_speed * (abs(speed - ideal_speed) / speed_range)) # Smooth driving bonus max_steering_variance = 0.1 # configured to the acceptable steering variance for full bonus smoothness_factor = np.std(action_list) reward_components[’smooth_driving_bonus’] = np.exp(-temp_smoothness * (smoothness_factor / max_steering_variance)) # Front distance bonus max_distance_view = 20 reward_components[’front_distance_bonus’] = np.clip(distance / max_distance_view, 0, 1) # Sum the individual rewards for the total reward total_reward = sum(reward_components.values()) # Ensure penalties take precedence if collision or speed < inactivity_threshold: total_reward = -1 # Apply max negative reward for collision or inactivity # Normalize the total reward to lie between -1 and 1 total_reward = np.clip(total_reward, -1, 1) return total_reward, reward_components
G.1.2 Humanoid Locomotion
In Box G.1.2, we present the best-performing reward from REvolve with an average velocity of $5.67$. In Box G.1.2, the best-performing reward from REvolve Auto has an average velocity of $4.71$.
Box 15: Best reward function from REvolve: Humanoid Locmotion def compute_reward(observation: np.array) -> (float, dict): # Constants and Parameters speed_temp = 0.1 orientation_temp = 0.05 health_temp = 1.0 # Extract relevant components from the observation speed = observation[22] # x-velocity component for reward z_position = observation[0] # z-coordinate of the torso orientation = observation[1:5] # orientation of the torso (x, y, z, w) # Reward Components speed_reward = np.exp(speed * speed_temp) - 1 # Encourage forward movement, exponentially orientation_penalty = -np.var(orientation) * orientation_temp # Penalize variance in orientation to encourage stability health_penalty = 0 if 1.0 < z_position < 2.0 else -np.exp(-abs(z_position - 1.5) * health_temp) # Penalty for unhealthy z-position # Total Reward Calculation total_reward = speed_reward + orientation_penalty + health_penalty reward_components = { ’speed_reward’: speed_reward, ’orientation_penalty’: orientation_penalty, ’health_penalty’: health_penalty } return total_reward, reward_components
Box 16: Best reward function from REvolve Auto: Humanoid Locomotion def compute_reward(observation: np.ndarray) -> (float, dict): # Constants and Parameters speed_reward_temp = 0.1 health_penalty_temp = 0.01 orientation_reward_temp = 0.1 health_threshold = 0.8 # Assumed threshold for the health component before penalties kick in more aggressively # Extracting necessary components from observation for computation z_position = observation[0] # Assuming first value represents torso z-position speed = observation[22] # Assuming this represents the forwards velocity # Speed Reward Component: Encourages forward movement speed_reward = speed * np.exp(speed_reward_temp * speed) # Health Penalty Component: Penalizes unhealthy states less aggressively is_unhealthy = 1 if z_position < 1.0 or z_position > 2.0 else 0 health_penalty = -np.exp(health_penalty_temp * is_unhealthy * (2.0 - z_position) if z_position < health_threshold else health_penalty_temp * is_unhealthy * 0.5) # Orientation Reward Component: Encourages maintaining upright torso orientation x_orientation, y_orientation = observation[1], observation[2] # Assumed orientation metrics orientation_reward = (1 - x_orientation**2 - y_orientation**2) * np.exp(orientation_reward_temp * (1 - x_orientation**2 - y_orientation**2)) # Total Reward Calculation total_reward = speed_reward + health_penalty + orientation_reward # Reward Components Dictionary reward_components = { ’speed_reward’: speed_reward, ’health_penalty’: health_penalty, ’orientation_reward’: orientation_reward } return total_reward, reward_components
G.1.3 Adroit Hand
”In Box G.1.3, we present the best-performing reward from REvolve with a fitness score of $0.81$. In Box G.1.3, the best-performing reward from REvolve Auto has a fitness score of $0.64$.”
Box 17: Best reward function from REvolve: Adroit Hand def compute_reward(observation, joint_velocities, joint_forces): # Example observables goal_pos = observation[28] # Using door_pos as an approximation for the goal position palm_to_handle = np.linalg.norm(observation[35:38]) # Magnitude of vector from palm to handle door_hinge_angle = observation[28] completion_threshold = 1.35 # radians, as per task success definition # Parameters dist_temp = 10.0 # Temperature parameter for distance cost transformation hinge_temp = 5.0 # Temperature parameter for hinge reward transformation # Components mutation: Distance cost modification with exponential decay dist_cost = -np.exp(-palm_to_handle / dist_temp) # Door hinge reward: reward for rotating the door handle door_hinge_reward = np.exp((door_hinge_angle - completion_threshold) / hinge_temp) if door_hinge_angle < completion_threshold else 1.0 # Checking if the door is sufficiently opened to consider the episode a success if door_hinge_angle >= completion_threshold: completion_bonus = 100.0 # Significant bonus for completing the task else: completion_bonus = 0.0 # Total reward calculation total_reward = dist_cost + door_hinge_reward + completion_bonus # Returning the total reward and the individual components as a dictionary return total_reward, {’dist_cost’: dist_cost, ’door_hinge_reward’: door_hinge_reward, ’completion_bonus’: completion_bonus} ~
Box 18: Best reward function from REvolve Auto: Adroit Hand def compute_reward(observation: np.ndarray) -> (float, dict): # Defining temperature parameters for the transformations temp_position = 0.1 # Temperature for position reward transformation temp_door_open = 0.05 # Temperature for door open reward transformation # Extract necessary components from the observation position_reward = observation[-2] # Assuming last but one value corresponds to positional info door_open_reward = observation[-1] # Assuming last value indicates door being open or not # Apply transformations with temperature parameters transformed_position_reward = np.exp(temp_position * position_reward) - 1 # Subtracting 1 to ensure a baseline of 0 transformed_door_open_reward = np.exp(temp_door_open * door_open_reward) - 1 # Calculate total reward total_reward = transformed_position_reward + transformed_door_open_reward # Constructing the dictionary of individual reward components reward_components = { ’position_reward’: transformed_position_reward, ’door_open_reward’: transformed_door_open_reward } return total_reward, reward_components